OpenStudy (anonymous):

solve the system by substitution -4.5x - 2y = -12.5 3.25x - y = -0.75

OpenStudy (anonymous):

Solve for a variable in either equation. The second one looks easiest to solve for y. 3.25x - y = -0.75 -y = -3.25x - 0.75 y = 3.25x + 0.75 Using this, we know what y equals. We can substitute this value in for y in the first equation. -4.5x - 2y = -12.5 -4.5x - 2(3.25x + 0.75) = -12.5 Now, you can solve for x using normal algebra. Once you do, you can use the equation you got for y earlier and plug in the x value. (y = 3.25x + 0.75)

OpenStudy (anonymous):

thank you